The average concentration of Carbon Monoxide in a city's air in 2006 was 4.5 ppm. The average atmospheric pressure of the city is 754 Torr.

(a) Calculate the partial pressure of CO.

(b) Calculate the number of CO molecules in 1.0L of this air at a temperature of 25°C
Answer #1

A) partial pressure of CO = mole fraction× total pressure

Partial pressure = 4.5×10-6×754 => 3.39×10-4 torr

B) n = 1×3.39×10-4/0.0821×298×760

n = 1.82×10-8 mol

Numner of molecules = 1.10×10-16